Tuesday, March 17th, 2009: Doug's Work Log

  • Marco is continuing to work on capacitive sensors calibration
  • A vacuum leak required Runa to fix a connector. He is trying to fill the Solar Tower with dry nitrogen (azoto) to lower the relative humidity so the temperature dependence of the mirror flat can be determined.
  • I made a plot of Acquisition Offsets for the field data collected on 16 March and sent to the AGW list. I encountered a problem with xmgrace and had to update port, sync macports and upgrade xorg-libXt.
  • Talked to Luca about the comments I made about FLAO-04. No major problems, just a few details.

-- DougMiller - 17 Mar 2009
Topic revision: r2 - 17 Mar 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback